According to Charles's law, for a fixed quantity of gas at constant pressure, which of the given quantities is constant?

VT

V×T

V+T


1
Expert's answer
2020-10-12T13:36:05-0400

Charles law states that volume of ideal gas is proportional to the absolute temperature of it.


i.e. volume of ideal gas = V unit.

and temperature = T K


Then according to the Charles law,

V ∝ T

Or, V = k × T

( k= proportionality constant)

Or, V/T = k

Or, V/T = constant



According to Charles's law, for a fixed quantity of gas at constant pressure V/T is constant.